A large balloon used to sample the upper atmosphere is filled with 525 m3 of hydrogen. What is the mass of the hydrogen (in kg)?
1
Expert's answer
2019-04-10T09:46:19-0400

We can find the mass of the hydrogen from the definition of the density:


"\rho = \dfrac{m}{V},"

here, "\rho = 0.089 kg/m^3" is the density of the hydrogen at standard temperature and pressure, "m" is the mass of the hydrogen and "V = 525 m^3" is the volume occupied by the hydrogen.

Then, from this formula we can find the mass of the hydrogen:


"m = \rho V = 0.089 \dfrac{kg}{m^3} \cdot 525 m^3 = 47 kg."

Answer:

"m = 47 kg."